OpenStudy (anonymous):

use quadratic formula to solve x^2=-6x-4

OpenStudy (solomonzelman):

\[x^2=-6x-4\]\[x^2\color{red} {+6x+4}=-6x-4\color{red} {+6x+4}\]\[x^2+6x+4=0\]\[\huge\color{blue}{x=\frac{-b±\sqrt{b^2-(4 \times a \times c)}}{2 \times a}}\] your a=1 b=6 c=4